# Kelvin waves in nonequilibrium universal dynamics of relativistic scalar field theories

### v1

@@ I Introduction
The concept of universality is a powerful framework for understanding complex many-body systems. Near equilibrium, it provides a way to associate phase transitions with universality classes based on a small set of system parameters, highlighting how vastly different systems can exhibit similar dynamical critical behaviour [1]. Beyond this, universality can also emerge far from equilibrium. Fluid turbulence represents one of the first scaling phenomena systematically discussed in physics [2]. Quantum coherence allows far-from-equilibrium scaling to emerge in low-energy as well as relativistic systems, in particular in the form of superfluid or quantum turbulence [3, 4, 5, 6, 7, 8, 9, 10, 11, 12, 13, 14] and different types of wave turbulence [15, 16, 17, 18, 19, 20, 21, 22], studied both in theory and experiment. Turbulence is commonly considered in an open-system setting, where driving and dissipation give rise to a stationary cascade that exhibits spatial scale invariance within an inertial range of energy scales.
@@
Universal dynamics far from equilibrium can also occur in closed systems exhibiting self-similar scaling in space and time. Such universality has been studied in analogy to (near-)equilibrium scaling, in connection with nonthermal fixed points [23, 24, 25, 26, 27, 28, 29, 30, 31], which give rise to nonequilibrium attractor solutions. This far-from-equilibrium universality has been experimentally observed in ultracold atomic gases in different dimensions and trapping geometries [9, 32, 33, 10, 34, 35, 21, 36, 37, 11, 38, 39, 40, 22, 12, 41, 42, 43, 44, 45, 46]. Many phenomena and properties of universal scaling dynamics have been theoretically predicted and studied in very different systems, ranging from ultracold gases over heavy-ion collisions to the early universe, [47, 48, 23, 24, 25, 49, 26, 50, 27, 51, 52, 53, 54, 55, 56, 57, 58, 59, 60, 61, 62, 63, 64, 65, 66, 28, 67, 68, 69, 70, 71, 72, 73, 74, 75, 29, 76, 77, 78, 79, 80, 81, 30, 82, 83, 84, 85, 86, 87, 88, 89, 90, 91, 92, 93, 94, 95, 96, 97, 98, 99, 100, 101, 102, 103, 104, 105, 106, 107, 108], for overviews cf. [109, 110, 111, 112, 113]. This includes studies of (wave) turbulence, as well as coarsening and phase-ordering kinetics [114, 115, 116]. The vast range of phenomena raises the question of what the relevant physics behind the apparent universality is, which universality classes exist and whether these classes can be divided into subclasses.
@@
The study of far-from-equilibrium universality in isolated systems has primarily focused on the properties of the time-dependent momentum distribution function, $f(t,p)$, which describes the occupancy of momentum modes over time. These functions reveal self-similar scaling behaviour near nonthermal fixed points, characterised by universal scaling exponents that are largely independent of the systems’ parameters or initial conditions. Relativistic scalar field theories with $\mathrm{O}(N)$ symmetry, for example, exhibit universal dynamics at low momenta according to $f(t,p)=t^{\,\alpha}f_{\mathrm{s}}(t^{\,\beta}p)$, with the exponents $\alpha=d\beta$ and $\beta=1/2$ [28]. The scaling function $f_{\mathrm{s}}$ is consistent across different values of $N$, whether in relativistic or nonrelativistic models, and the underlying physics is connected to the self-similar transport of particle numbers toward lower momenta and the growth of a quasi condensate in the long-wavelength modes below a characteristic scale $p_{\mathrm{IR}}(t)\sim t^{-\beta}$.
@@
Large-$N$ kinetic theories provide a successful description of this scaling dynamics and the exponents, which can be traced to weak elastic collisions of Goldstone quasiparticles with free dispersion and momentum-dependent effective quartic coupling [23, 24, 25, 117, 28, 60, 118, 29, 30, 81, 106, 108], representing the excitations of the phase-angle differences of the different components [81, 31]. For small values of $N$, however, where the excitations of the total field amplitude become relevant, the precise description of the universality remains unsatisfactory, and alternative mechanisms, such as coarsening dynamics of topological defects are anticipated to play a significant role [119, 120, 26, 27, 54, 52, 59, 74, 76, 11, 86, 87, 98, 42, 121]. Recent studies indicate that for systems with a small number of field components, distinct initial conditions may lead to different scaling exponents [74, 42, 97, 103], where distinctly sub-diffusive scaling, $\beta\ll 1/2$, has been associated with specific defect-driven dynamics, e.g., $\beta\approx 1/5$ for vortices in two spatial dimensions [74, 11, 88, 102].
@@
For the relativistic $\mathrm{O}(N)$-symmetric nonlinear Klein-Gordon vector models in both two and three spatial dimensions, a number of different, $N$-dependent topological defects are expected to contribute to the dynamics in the infrared [61, 76, 102], which includes vortex defects in $\mathrm{O}(1)$. However, for $\mathrm{O}(N)$ theories in three (spatial) dimensions, including the case $\mathrm{O}(1)$, the self-similar scaling evolution of the distribution function has been consistent with that of large-$N$ descriptions and in agreement with nonrelativistic $\mathrm{U}(N)$-symmetric nonlinear Schrödinger models  [25, 28, 61, 29, 30, 81, 122, 31]. This has been corroborated by observing the scaling behaviour of $f(t,p)$ based on classical-statistical (Truncated-Wigner) real-time lattice simulations. Nonetheless, unequal-time correlation functions, which also give information about the nonequilibrium excitation spectrum, have revealed markedly different dominant excitation peaks for $N=1,2$ compared to $N\geq 3$, where the large-$N$ excitation indeed dominates [31].
@@
Here we study the underlying mechanisms of the self-similar infrared transport for the relativistic single-component $\mathrm{O}(1)$ theory in two and three spatial dimensions using unequal-time two-point correlation functions. By providing information on both occupancies and dispersion relations, the latter observables allow us to distinguish the different excitations contributing to the scaling and also to ascertain which excitation is dominant. Based on numerical simulations in three spatial dimensions, the dominant excitation in such correlators at large $N$ has been identified as the one with the (complex) quasiparticle dispersion [123, 31] that agrees with the kinetic theory picture. An excitation similar to the large-$N$ excitation has also been found for the relativistic $N=2$ case at low momenta [31], whereas it is absent for $N=1$. On the other hand, a very specific, so-called “transport peak” has been identified for both nonrelativistic $\mathrm{U}(1)$ and in the nonrelativistic limit of relativistic $\mathrm{O}(1)$ models [122, 31], with the same momentum and time dependence of the dispersion relation, as well as damping rate, while being distinct from the large-$N$ excitation. This peak appears to be present for the cases of $\mathrm{O}(2)$ and $\mathrm{O}(3)$ as well, however, it distinctly dominates only the $N=1$ theory [31]. The origin of this transport peak has not been understood yet and is not predicted by any current kinetic or low-energy effective description of nonthermal fixed points.
@@
In anticipation of its potential connection to topological defects, in this work, we also extract and analyse the vortex dynamics for the $\mathrm{O}(1)$ theory in both two and three dimensions using topological defect data extracted from lattice configurations. In the context of nonthermal fixed points, two-dimensional relativistic theories have not been extensively studied yet [76, 120, 54]. However, they do provide an interesting toy model since, intuitively speaking, topological effects are more important in lower dimensions and for lower $N$. By carefully investigating the transport peak, we find a direct connection between the self-similar infrared transport near the nonthermal fixed point and vortex line defects in three dimensions in terms of Kelvin waves [124, 125, 126, 127], which arise as quantised excitations on vortices present in scalar theories, and which are well-investigated in the context of vortex dynamics and quantum turbulence [128, 129, 17, 130, 131, 132, 133, 19, 134, 135, 20, 136]. In two dimensions, we observe a similar transport peak, where the underlying physics is connected to the two-dimensional analogues of Kelvin waves, which arise as vortex-displacement excitations [137, 138, 139]. While the general picture of nonlinear wave propagation in the strongly correlated regime near a nonthermal fixed point still applies for both small and large values of $N$, with the present study, we distinctly identify the relevant degrees of freedom in these theories, highlighting the importance of topological defects for a single-component scalar field theory.
@@ II Model and observables
where $\int_{t,\mathbb{x}}\equiv\int\mathrm{d}t\int\mathrm{d}^{d}x$, summation over repeated indices is implied, $m$ is the bare mass and $\lambda$ is the coupling constant. In the weak-coupling regime $\lambda\ll 1$ considered in this work, the quantum dynamics can be accurately mapped onto classical-statistical field theory (Truncated-Wigner approximation) for highly occupied systems at not too late times [140]. In classical-statistical simulations, one samples the fields over initial conditions,
@@
In the numerical approach, we discretise $\phi(t,\mathbf{x})$ and $\pi(t,\mathbf{x})$ on a spatial lattice with spacing $a_{s}$, volume $V=(N_{s}a_{s})^{d}$ and time step $a_{t}=0.05\,a_{s}$. In $d=2$ dimensions, we use $N_{s}=2048$, $a_{s}=0.1\,Q^{-1}$, $m/Q=2$, $n_{0}=80$. In $d=3$ dimensions, we use $N_{s}=256$, $a_{s}=0.8\,Q^{-1}$, $m/Q=0.25$, $n_{0}=100$, unless stated otherwise. Any dependence on the coupling constant drops out after the classical equations of motion and the initial conditions are rescaled according to $\phi\rightarrow\lambda^{1/2}\phi$, $\pi\rightarrow\lambda^{1/2}\pi$. This property stems from the underlying classical limit $\hbar\to 0$ or, equivalently, the weak coupling limit $\lambda\rightarrow 0$ while $\lambda f$ remains unchanged. The figures shown in the following sections all assume quantities in units of $Q$.
@@
For the computation of the spectral function, the linear-response approach of [31] similar to [141, 122] is employed. This involves perturbing the system, at time $t$, as $\phi\rightarrow\phi+\delta\phi$ with a source term $j(t,\mathbf{x})=j^{0}(\mathbf{x})\delta\left(t-t^{\prime}\right)$, drawn from a random Gaussian distribution that satisfies $\left\langle j^{0}(\mathbf{p})\left(j^{0}(\mathbf{q})\right)^{*}\right\rangle_{j}=V\delta_{\mathbf{p},\mathbf{q}}$. The perturbation $\delta\phi$ is time evolved according to the linearised equations of motion $\partial_{t}^{\prime}\delta\phi=\delta\pi$ and
@@
where $t\equiv\left(t^{\prime}+t^{\prime\prime}\right)/2$ is the (central) time. In practice, we evaluate the correlations at $t^{\prime\prime}=t$. This approximation can be justified provided the correlations do not change significantly over times of the order of their oscillation periods along the relative time $\Delta t$, which is fulfilled during the self-similar scaling dynamics [122]. Therefore, in our simulations, we compute the Fourier transforms while holding $t$ fixed, as
@@
and zero padding to smoothen the resulting curves, and have checked that these do not alter the form of the displayed peaks. Expectation values of observables are obtained by averaging over the field trajectories that are solutions to the classical equations of motion. The results shown in the following use correlation functions averaged over 100 realisations in two and 40 realisations in three dimensions. Real-time classical-statistical simulations have been extensively used to study the dynamics of isolated systems in corresponding regimes of applicability for similar initial conditions [31, 102, 142, 23, 120, 143, 144, 59, 68, 61, 118, 83, 26, 27, 52, 67, 74, 69, 76, 86, 87, 93, 96, 98, 104, 103].

### v2

@@ I Introduction
The concept of universality is a powerful framework for understanding complex many-body systems. Near equilibrium, it provides a way to associate phase transitions with universality classes based on a small set of system parameters, highlighting how vastly different systems can exhibit similar dynamical critical behaviour Hohenberg and Halperin (1977). Beyond this, universality can also emerge far from equilibrium. Fluid turbulence represents one of the first scaling phenomena systematically discussed in physics Frisch (1995). Quantum coherence allows far-from-equilibrium scaling to emerge in low-energy as well as relativistic systems, in particular in the form of superfluid or quantum turbulence Svistunov (2001); Volovik (2004); Vinen (2006); Tsubota (2008); Proment et al. (2011); Micha and Tkachev (2003); Henn et al. (2009); Kwon et al. (2014); Johnstone et al. (2019); Glidden et al. (2021); Müller et al. (2021); Müller and Krstulovic (2024) and different types of wave turbulence Zakharov et al. (1992); Nazarenko (2011); Kozik and Svistunov (2004); Mueller et al. (2007); L’vov and Nazarenko (2010); Boué et al. (2011); Navon et al. (2016, 2019), studied both in theory and experiment. Turbulence is commonly considered in an open-system setting, where driving and dissipation give rise to a stationary cascade that exhibits spatial scale invariance within an inertial range of energy scales.
@@
Universal dynamics far from equilibrium can also occur in closed systems exhibiting self-similar scaling in space and time. Such universality has been studied in analogy to (near-)equilibrium scaling, in connection with nonthermal fixed points Berges et al. (2008); Berges and Hoffmeister (2009); Scheppach et al. (2010); Nowak et al. (2012); Schole et al. (2012); Piñeiro Orioli et al. (2015); Walz et al. (2018); Chantesana et al. (2019); Boguslavski and Piñeiro Orioli (2020), which give rise to nonequilibrium attractor solutions. This far-from-equilibrium universality has been experimentally observed in ultracold atomic gases in different dimensions and trapping geometries Henn et al. (2009); Gring et al. (2012); Smith et al. (2013); Kwon et al. (2014); Langen et al. (2015); Navon et al. (2015, 2016); Rauer et al. (2018); Gauthier et al. (2019); Johnstone et al. (2019); Eigen et al. (2018); Prüfer et al. (2018); Erne et al. (2018); Navon et al. (2019); Glidden et al. (2021); Garc\́mathrm{i}a-Orozco et al. (2022); Lannig et al. (2023); Martirosyan et al. (2024a); Gazo et al. (2023); Moreno-Armijos et al. (2025); Martirosyan et al. (2024b). Many phenomena and properties of universal scaling dynamics have been theoretically predicted and studied in very different systems, ranging from ultracold gases over heavy-ion collisions to the early universe, Kodama et al. (2005); Mukerjee et al. (2007); Berges et al. (2008); Berges and Hoffmeister (2009); Scheppach et al. (2010); Barnett et al. (2011); Nowak et al. (2012, 2014); Schole et al. (2012); Berges and Sexty (2012); Karl et al. (2013, 2013); Gasenzer et al. (2014); Berges et al. (2014a); Marcuzzi et al. (2013); Langen et al. (2013); Berges et al. (2015a); Ewerz et al. (2015); Berges et al. (2015b); Moore (2016); Nessi et al. (2014); Gagel et al. (2014); Bertini et al. (2015); Babadi et al. (2015); Buchhold et al. (2016); Piñeiro Orioli et al. (2015); Hofmann et al. (2014); Maraga et al. (2015); Williamson and Blakie (2016a, b); Villois et al. (2016); Bourges and Blakie (2017); Chiocchetta et al. (2016); Karl and Gasenzer (2017); Berges et al. (2017); Walz et al. (2018); Deng et al. (2018); Mazeliauskas and Berges (2019); Baggaley and Barenghi (2018); Bland et al. (2018); Schmied et al. (2019a); Mikheev et al. (2019); Chantesana et al. (2019); Schlichting and Teaney (2019); Boguslavski et al. (2019); Berges et al. (2020); Williamson and Blakie (2019); Schmied et al. (2019b, c); Spitz et al. (2021); Gao et al. (2020); Groszek et al. (2020, 2021); Wheeler et al. (2021); Gresista et al. (2022); Liu et al. (2023); Mikheev et al. (2022); Heinen et al. (2022, 2023); Siovitz et al. (2023); Berges et al. (2024); Spitz et al. (2023); Heller et al. (2024); Noel and Spitz (2024); Siovitz et al. (2024); Spitz et al. (2024); Rosenhaus and Smolkin (2024); Rosenhaus and Falkovich (2024); Rosenhaus and Schubring (2024); Rosenhaus and Falkovich (2025), for overviews cf. Nowak et al. (2016); Berges (2016); Schmied et al. (2019d); Berges et al. (2021); Siovitz et al. . This includes studies of (wave) turbulence, as well as coarsening and phase-ordering kinetics Bray (1994); Puri and Wadhawan (2009); Cugliandolo (2015). The vast range of phenomena raises the question of what the relevant physics behind the apparent universality is, which universality classes exist and whether these classes can be divided into subclasses.
@@
The study of far-from-equilibrium universality in isolated systems has primarily focused on the properties of the time-dependent momentum distribution function, $f(t,p)$, which describes the occupancy of momentum modes over time. These functions reveal self-similar scaling behaviour near nonthermal fixed points, characterised by universal scaling exponents that are largely independent of the systems’ parameters or initial conditions. Relativistic scalar field theories with $\mathrm{O}(N)$ symmetry, for example, exhibit universal dynamics at low momenta according to $f(t,p)=t^{\,\alpha}f_{\mathrm{s}}(t^{\,\beta}p)$, with the exponents $\alpha=d\beta$ and $\beta=1/2$ Piñeiro Orioli et al. (2015). The scaling function $f_{\mathrm{s}}$ is consistent across different values of $N$, whether in relativistic or nonrelativistic models, and the underlying physics is connected to the self-similar transport of particle numbers toward lower momenta and the growth of a quasi condensate in the long-wavelength modes below a characteristic scale $p_{\mathrm{IR}}(t)\sim t^{-\beta}$.
@@
Large-NN kinetic theories provide a successful description of this scaling dynamics and the exponents, which can be traced to weak elastic collisions of Goldstone quasiparticles with free dispersion and momentum-dependent effective quartic coupling Berges et al. (2008); Berges and Hoffmeister (2009); Scheppach et al. (2010); Berges and Sexty (2011); Piñeiro Orioli et al. (2015); Berges et al. (2015b); Berges and Wallisch (2017); Walz et al. (2018); Chantesana et al. (2019); Mikheev et al. (2019); Rosenhaus and Falkovich (2024, 2025), representing the excitations of the phase-angle differences of the different components Mikheev et al. (2019); Boguslavski and Piñeiro Orioli (2020). For small values of $N$, however, where the excitations of the total field amplitude become relevant, the precise description of the universality remains unsatisfactory, and alternative mechanisms, such as coarsening dynamics of topological defects are anticipated to play a significant role Nowak et al. (2011); Gasenzer et al. (2012); Nowak et al. (2012); Schole et al. (2012); Gasenzer et al. (2014); Karl et al. (2013); Ewerz et al. (2015); Karl and Gasenzer (2017); Deng et al. (2018); Johnstone et al. (2019); Schmied et al. (2019b, c); Siovitz et al. (2023); Lannig et al. (2023); Huh et al. (2023). Recent studies indicate that for systems with a small number of field components, distinct initial conditions may lead to different scaling exponents Karl and Gasenzer (2017); Lannig et al. (2023); Heinen et al. (2023); Siovitz et al. (2024), where distinctly sub-diffusive scaling, $\beta\ll 1/2$, has been associated with specific defect-driven dynamics, e.g., $\beta\approx 1/5$ for vortices in two spatial dimensions Karl and Gasenzer (2017); Johnstone et al. (2019); Spitz et al. (2021); Noel and Spitz (2024).
@@
For the relativistic $\mathrm{O}(N)$-symmetric nonlinear Klein-Gordon vector models in both two and three spatial dimensions, a number of different, $N$-dependent topological defects are expected to contribute to the dynamics in the infrared Moore (2016); Deng et al. (2018); Noel and Spitz (2024), which includes vortex defects in $\mathrm{O}(1)$. The topological connection between the $\mathrm{O}(1)$ and the nonrelativistic $\mathrm{U}(1)$ theory, and hence, the existence and definition of vortices in the relativistic $\mathrm{O}(1)$ model originates from an emerging mass gap. Therefore, at low energies, the field and its canonical conjugate are mapped to the real and imaginary parts of the complex $\mathrm{U}(1)$ field. However, for $\mathrm{O}(N)$ theories in three (spatial) dimensions, including the case $\mathrm{O}(1)$, the self-similar scaling evolution of the distribution function has been consistent with that of large-$N$ descriptions and in agreement with nonrelativistic $\mathrm{U}(N)$-symmetric nonlinear Schrödinger models  Scheppach et al. (2010); Piñeiro Orioli et al. (2015); Moore (2016); Walz et al. (2018); Chantesana et al. (2019); Mikheev et al. (2019); Piñeiro Orioli and Berges (2019); Boguslavski and Piñeiro Orioli (2020). This has been corroborated by observing the scaling behaviour of $f(t,p)$ based on classical-statistical (Truncated-Wigner) real-time lattice simulations. Nonetheless, unequal-time correlation functions, which also give information about the nonequilibrium excitation spectrum, have revealed markedly different dominant excitation peaks for $N=1,2$ compared to $N\geq 3$, where the large-$N$ excitation indeed dominates Boguslavski and Piñeiro Orioli (2020).
@@
Here we study the underlying mechanisms of the self-similar infrared transport for the relativistic single-component $\mathrm{O}(1)$ theory in two and three spatial dimensions using unequal-time two-point correlation functions. By providing information on both occupancies and dispersion relations, the latter observables allow us to distinguish the different excitations contributing to the scaling and also to ascertain which excitation is dominant. Based on numerical simulations in three spatial dimensions, the dominant excitation in such correlators at large $N$ has been identified as the one with the (complex) quasiparticle dispersion Shen and Berges (2020); Boguslavski and Piñeiro Orioli (2020) that agrees with the kinetic theory picture. An excitation similar to the large-$N$ excitation has also been found for the relativistic $N=2$ case at low momenta Boguslavski and Piñeiro Orioli (2020), whereas it is absent for $N=1$. On the other hand, a very specific, so-called “transport peak” has been identified for both nonrelativistic $\mathrm{U}(1)$ and in the nonrelativistic limit of relativistic $\mathrm{O}(1)$ models Piñeiro Orioli and Berges (2019); Boguslavski and Piñeiro Orioli (2020), with the same momentum and time dependence of the dispersion relation, as well as damping rate, while being distinct from the large-$N$ excitation. This peak appears to be present for the cases of $\mathrm{O}(2)$ and $\mathrm{O}(3)$ as well, however, it distinctly dominates only the $N=1$ theory Boguslavski and Piñeiro Orioli (2020). The origin of this transport peak has not been understood yet and is not predicted by any current kinetic or low-energy effective description of nonthermal fixed points.
@@
In anticipation of its potential connection to topological defects, in this work, we also extract and analyse the vortex dynamics for the $\mathrm{O}(1)$ theory in both two and three dimensions using topological defect data extracted from lattice configurations. In the context of nonthermal fixed points, two-dimensional relativistic theories have not been extensively studied yet Deng et al. (2018); Gasenzer et al. (2012, 2014). However, they do provide an interesting toy model since, intuitively speaking, topological effects are more important in lower dimensions and for lower $N$. By carefully investigating the transport peak, we find a direct connection between the self-similar infrared transport near the nonthermal fixed point and vortex line defects in three dimensions in terms of Kelvin waves Thomson (Lord Kelvin); Pitaevskii (1961); Gross (1961); Donnelly (2005), which arise as quantised excitations on vortices present in scalar theories, and which are well-investigated in the context of vortex dynamics and quantum turbulence Vinen (2001); Vinen et al. (2003); Kozik and Svistunov (2004, 2005a, 2005b, 2008, 2009); L’vov and Nazarenko (2010); Krstulovic (2012); Laurie et al. (2010); Boué et al. (2011, 2015). In two dimensions, we observe a similar transport peak, where the underlying physics is connected to the two-dimensional analogues of Kelvin waves, which arise as vortex-displacement excitations Isoshima and Machida (1999); Simula (2013, 2018). While the general picture of nonlinear wave propagation in the strongly correlated regime near a nonthermal fixed point still applies for both small and large values of $N$, with the present study, we distinctly identify the relevant degrees of freedom in these theories, highlighting the importance of topological defects for a single-component scalar field theory.
@@ II Model and observables
where $\int_{t,\mathbb{x}}\equiv\int\mathrm{d}t\int\mathrm{d}^{d}x$, summation over repeated indices is implied, $m$ is the bare mass and $\lambda$ is the coupling constant. In the weak-coupling regime $\lambda\ll 1$ considered in this work, the quantum dynamics can be accurately mapped onto classical-statistical field theory (Truncated-Wigner approximation) for highly occupied systems at not too late times Polkovnikov (2010). In classical-statistical simulations, one samples the fields over initial conditions,
@@
In the numerical approach, we discretise $\phi(t,\mathbf{x})$ and $\pi(t,\mathbf{x})$ on a spatial lattice with spacing $a_{s}$, volume $V=(N_{s}a_{s})^{d}$ and time step $a_{t}=0.05\,a_{s}$. In $d=2$ dimensions, we use $N_{s}=2048$, $a_{s}=0.1\,Q^{-1}$, $m/Q=2$, $n_{0}=80$. In $d=3$ dimensions, we use $N_{s}=256$, $a_{s}=0.8\,Q^{-1}$, $m/Q=0.25$, $n_{0}=100$, unless stated otherwise. We use a fourth-order accurate finite-difference scheme for the second-order spatial derivatives as described in Micha and Tkachev (2003). The time evolution is performed using a leapfrog algorithm in propagating Eqs. (LABEL:eq:class_EOM), which conserves the energy density and where the fields $\phi$ and $\pi$ are half a timestep apart. We furthermore impose periodic boundary conditions for the spatial lattice.
Any dependence on the coupling constant drops out after the classical equations of motion and the initial conditions are rescaled according to $\phi\rightarrow\lambda^{1/2}\phi$, $\pi\rightarrow\lambda^{1/2}\pi$. This property stems from the underlying classical limit $\hbar\to 0$ or, equivalently, the weak coupling limit $\lambda\rightarrow 0$ while $\lambda f$ remains unchanged. The figures shown in the following sections all assume quantities in units of $Q$.
@@
For the computation of the spectral function, the linear-response approach of Boguslavski and Piñeiro Orioli (2020) similar to Boguslavski et al. (2018); Piñeiro Orioli and Berges (2019) is employed. This involves perturbing the system, at time $t$, as $\phi\rightarrow\phi+\delta\phi$ with a source term $j(t,\mathbf{x})=j^{0}(\mathbf{x})\delta\left(t-t^{\prime}\right)$, drawn from a random Gaussian distribution that satisfies $\left\langle j^{0}(\mathbf{p})\left(j^{0}(\mathbf{q})\right)^{*}\right\rangle_{j}=V\delta_{\mathbf{p},\mathbf{q}}$. The perturbation $\delta\phi$ is time evolved according to the linearised equations of motion $\partial_{t}^{\prime}\delta\phi=\delta\pi$ and
@@
where $t\equiv\left(t^{\prime}+t^{\prime\prime}\right)/2$ is the (central) time. In practice, we evaluate the correlations at $t^{\prime\prime}=t$. This approximation can be justified provided the correlations do not change significantly over times of the order of their oscillation periods along the relative time $\Delta t$, which is fulfilled during the self-similar scaling dynamics Piñeiro Orioli and Berges (2019). Therefore, in our simulations, we compute the Fourier transforms while holding $t$ fixed, as
@@
and zero padding to smoothen the resulting curves, and have checked that these do not alter the form of the displayed peaks. Expectation values of observables are obtained by averaging over the field trajectories that are solutions to the classical equations of motion. The results shown in the following use correlation functions averaged over 100 realisations in two and 40 realisations in three dimensions. Real-time classical-statistical simulations have been extensively used to study the dynamics of isolated systems in corresponding regimes of applicability for similar initial conditions Boguslavski and Piñeiro Orioli (2020); Noel and Spitz (2024); Mikheev et al. (2025); Berges et al. (2008); Gasenzer et al. (2012); Berges et al. (2014b, c); Ewerz et al. (2015); Maraga et al. (2015); Moore (2016); Berges and Wallisch (2017); Boguslavski et al. (2019); Nowak et al. (2012); Schole et al. (2012); Karl et al. (2013); Hofmann et al. (2014); Karl and Gasenzer (2017); Williamson and Blakie (2016a); Deng et al. (2018); Schmied et al. (2019b, c); Gresista et al. (2022); Heinen et al. (2022); Siovitz et al. (2023); Spitz et al. (2024); Siovitz et al. (2024).